How many grams of berylium chloride are needed to make 125 mL of a .050 M solution? And how do I get the answer?
1
Expert's answer
2012-02-16T10:55:54-0500
0.05М = 0.05 moles in 1000 ml,so in 125 ml will be 125*0.05/1000=0.00625
moles
n=m/MW
m=n*MW=0.00625*80=0.5 g